A company currently pays a dividend of $2.4 per share (D0 = $2.4). It is estimated that the company's dividend will grow at a rate of 22% per year for the next 2 years, and then at a constant rate of 7% thereafter. The company's stock has a beta of 1.5, the risk-free rate is 8%, and the market risk premium is 3%. What is your estimate of the stock's current price? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Required return=risk free rate+Beta*market risk premium

=8+(1.5*3)=12.5%

D1=(2.4*1.22)=2.928

D2=(2.928*1.22)=3.57216

Value after year 2=(D2*Growth rate)/(Required return-Growth rate)

=(3.57216*1.07)/(0.125-0.07)

=69.4947491

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=2.928/1.125+3.57216/1.125^2+69.4947491/1.125^2

=$60.33(Approx).